OpenStudy (anonymous):

factor the polynomial. 4. 12x^5 – 20x^4 + 16x^2

OpenStudy (brinazarski):

First, you should find something that they're all divisible by :)

OpenStudy (anonymous):

I can't go past 4x^293x^3-5x^2+4)

OpenStudy (brinazarski):

Woah what?

OpenStudy (anonymous):

4x^2(3x^3-5x^2+4)

OpenStudy (anonymous):

12x^5 - 20x^4 + 16x^2 = 4*3x^2 x^3 - 4*5 x^2 x^2 + 4*4 x^2 = 4x^2(3x^3 - 5x^2 + 4)

OpenStudy (anonymous):

fine?

OpenStudy (anonymous):

yeah no it was exactly right fontez